OpenStudy (anonymous):

if f(x)=x^2-3 , then f(a-b) what does that equal

OpenStudy (amistre64):

f(x) defines an equation in terms of "x"; whih you have done. f(9) would mean to use "9" as the value of "x" to solve the equation. What they have given you is: f(a-b) which simply means that you use "a-b" in place of x and solve.

OpenStudy (amistre64):

f(x) = x^2 -3 f(a-b) = (a-b)^2 -3

OpenStudy (amistre64):

f(a-b) = a^2 -2ab +b^2 -3 and thats all you can do with it. Does that make sense?

OpenStudy (anonymous):

yessss thank you so muchh !!

OpenStudy (anonymous):

by any chance do you know this one: if f(x)= x-2 / x+1 , then f(n+1) equals what?